# One Backpropagation in Two Tower Recommendation Models

### v1

@@ 2. Two Tower Recommendation Models
$\mathbf{u}_{i}=UserEncoder(u_{i}),$
$\mathbf{v}_{j}=ItemEncoder(v_{j}).$
@@ 5.1. Experiment Settings
Evaluation metrics: We adopt the mostly used performance metrics in the OCCF problem for personalized top-$K$ recommendation, including Precision (P), Recall (R), F1-Score (F1), and Normalized Discounted Cumulative Gain (NDCG). For their wide usage, we do not detail their calculations here. We use them to evaluate the top-$5$, top-$10$, and top-$20$ recommendation performance.
Parameter Settings: To make fair comparisons, we use the same number of negative samples for all algorithms, and use the grid search to fine-tune the hyperparameters of each algorithm for its best Precision performance. More detailed hyperparameter settings are documented in the README file of the code. The codes is implemented with Pytorch  [^4] Codes are released at: https://anonymous.4open.science/r/OneBP. Computations were conducted on a cloud server equipped with 32 Intel Xeon Processor CPU(Skylake, IBRS), Tesla V100 GPU*1 (36GB).
@@ 5.2. Overall results
Table 3 compares the computation time per epoch on the four datasets, where our OneBP requires the least ones. The savings of computation time become more significant in larger datasets. This is not unexpected. As our solution does not require advanced neural encoding, sophisticated negative sampling and complicated loss computing, its computation complexity for model training can be greatly reduced. Further computation savings can be attributed to that our OneBP does not need to compute loss gradients for updating users’ encodings.
@@
Type Method P@5 R@5 F1@5 NDCG@5 P@10 R@10 F1@10 NDCG@10 P@20 R@20 F1@20 NDCG@20
MovieLens-100k
Encoder MF 0.4155 0.1434 0.1850 0.4430 0.3496 0.2308 0.2328 0.4181 0.2824 0.3529 0.2606 0.4183
LightGCN 0.4263 0.1458 0.1881 0.4531 0.3587 0.2313 0.2351 0.4269 0.2877 0.3474 0.2609 0.4236
NGCF 0.4176 0.1452 0.1871 0.4469 0.3582 0.2356 0.238 0.4264 0.2875 0.3597 0.2652 0.4248
Negative DNS 0.4214 0.1460 0.1882 0.4501 0.3531 0.2317 0.2342 0.4226 0.2836 0.3544 0.2614 0.4228
PNS 0.3979 0.1338 0.1742 0.4231 0.3415 0.218 0.223 0.4026 0.2741 0.3373 0.2511 0.4016
HarS 0.4227 0.1455 0.1878 0.4516 0.3527 0.2312 0.2342 0.4239 0.2852 0.3557 0.2624 0.4247
2stS 0.3718 0.1217 0.1597 0.3929 0.3188 0.2027 0.2085 0.3715 0.2574 0.3184 0.2365 0.3717
Loss CML 0.3796 0.1174 0.1562 0.3963 0.3314 0.2022 0.2109 0.3793 0.272 0.3243 0.2458 0.3811
SFCML 0.4174 0.1432 0.1846 0.4429 0.3544 0.2292 0.2321 0.4202 0.2814 0.3389 0.2542 0.4147
DPCML 0.3869 0.1337 0.1717 0.4115 0.3207 0.2108 0.213 0.3836 0.2647 0.3287 0.2435 0.3833
DCL 0.4291 0.1501 0.1928 0.4587 0.3608 0.2384 0.2402 0.4319 0.2845 0.3542 0.2621 0.4257
HCL 0.4263 0.1499 0.1924 0.4576 0.3576 0.2365 0.2389 0.4295 0.2814 0.3513 0.2603 0.4229
BUIR 0.4146 0.1387 0.1804 0.4373 0.3491 0.2258 0.2297 0.4121 0.2743 0.3445 0.2529 0.407
Proposed OneBP 0.453 0.1559 0.2015 0.4822 0.3811 0.2494 0.2529 0.4535 0.3032 0.373 0.2776 0.4501
5.57$\uparrow$ 3.86$\uparrow$ 4.51$\uparrow$ 5.12$\uparrow$ 5.63$\uparrow$ 4.61$\uparrow$ 5.29$\uparrow$ 5.00$\uparrow$ 5.39$\uparrow$ 3.70$\uparrow$ 4.68$\uparrow$ 5.73$\uparrow$
MovieLens-1M
Encoder MF 0.4046 0.0962 0.1355 0.4234 0.3510 0.1584 0.1816 0.3930 0.2922 0.2477 0.2166 0.3780
LightGCN 0.4121 0.0955 0.1350 0.4327 0.3537 0.1584 0.1792 0.3981 0.2921 0.2401 0.2132 0.3791
NGCF 0.4192 0.1014 0.1422 0.4397 0.3608 0.1643 0.1879 0.4059 0.2983 0.2560 0.2234 0.3888
Negative DNS 0.4080 0.0975 0.1372 0.4276 0.3558 0.1606 0.1841 0.3980 0.2948 0.2504 0.2190 0.3819
PNS 0.3931 0.0950 0.1334 0.4129 0.3362 0.1542 0.1759 0.3802 0.2773 0.2386 0.2079 0.3639
HarS 0.3453 0.0719 0.1042 0.3687 0.2770 0.1086 0.1293 0.3201 0.2049 0.1544 0.1424 0.2781
2stS 0.3772 0.0930 0.1295 0.3978 0.3261 0.1507 0.1704 0.3679 0.2689 0.2362 0.2029 0.3521
Loss CML 0.3425 0.0753 0.1077 0.3554 0.3047 0.1277 0.1497 0.3341 0.258 0.2055 0.1847 0.3226
SFCML 0.3968 0.0959 0.1348 0.4136 0.3481 0.1573 0.1808 0.3875 0.2912 0.243 0.2154 0.3741
DPCML 0.3883 0.088 0.125 0.4079 0.3359 0.1424 0.1665 0.3755 0.2758 0.2198 0.1983 0.3538
DCL 0.4133 0.0953 0.1350 0.4355 0.3512 0.152 0.177 0.3967 0.2867 0.2328 0.2084 0.3732
HCL 0.4125 0.0957 0.1355 0.4345 0.3515 0.1528 0.1776 0.3972 0.2867 0.2336 0.2086 0.3736
BUIR 0.3875 0.0942 0.1318 0.4089 0.3314 0.152 0.1728 0.376 0.2714 0.2332 0.2023 0.3581
Proposed OneBP 0.4300 0.1071 0.1487 0.4512 0.3709 0.1746 0.1966 0.4182 0.3069 0.2690 0.2316 0.4036
2.58$\uparrow$ 5.62$\uparrow$ 4.57$\uparrow$ 2.62$\uparrow$ 2.80$\uparrow$ 6.27$\uparrow$ 4.63$\uparrow$ 3.03$\uparrow$ 2.88$\uparrow$ 5.08$\uparrow$ 3.67$\uparrow$ 3.81$\uparrow$
Gowalla
Encoder MF 0.0739 0.0757 0.0621 0.1016 0.0560 0.1122 0.0623 0.1076 0.0422 0.1650 0.0587 0.1230
LightGCN 0.0845 0.0888 0.0725 0.1166 0.0631 0.1297 0.0713 0.1232 0.0465 0.1864 0.0642 0.1394
NGCF 0.0768 0.0808 0.0661 0.1051 0.0585 0.121 0.0664 0.1127 0.044 0.1769 0.0608 0.1291
Negative DNS 0.0737 0.0779 0.0633 0.1013 0.0563 0.1162 0.0636 0.1085 0.0421 0.1696 0.0580 0.1240
PNS 0.0586 0.0611 0.0502 0.0811 0.0439 0.0894 0.0494 0.0855 0.0323 0.1298 0.0445 0.0968
HarS 0.0719 0.0742 0.0609 0.0989 0.0544 0.1098 0.0607 0.1048 0.0411 0.16 0.0558 0.1189
2stS 0.0609 0.0599 0.0501 0.0832 0.0461 0.0883 0.0502 0.0871 0.0345 0.1285 0.0462 0.0982
Loss CML 0.0501 0.0603 0.0472 0.0705 0.0384 0.0901 0.0466 0.0782 0.0284 0.1285 0.0412 0.0902
SFCML 0.0899 0.0931 0.0766 0.1243 0.0672 0.1365 0.0756 0.1278 0.0496 0.1963 0.0689 0.1477
DPCML 0.0782 0.0759 0.0642 0.1042 0.059 0.1124 0.0645 0.1092 0.0432 0.1632 0.0584 0.1225
DCL 0.0761 0.0776 0.0642 0.1036 0.0575 0.1149 0.064 0.1093 0.0424 0.1651 0.0578 0.1236
HCL 0.0873 0.0885 0.0731 0.118 0.0653 0.1286 0.0723 0.1236 0.0477 0.1846 0.0649 0.1392
BUIR 0.0854 0.0872 0.072 0.1166 0.0651 0.1306 0.0725 0.1241 0.0484 0.1898 0.0661 0.1411
Proposed OneBP 0.0948 0.0971 0.0803 0.1282 0.0715 0.1431 0.0798 0.1356 0.0527 0.2059 0.0722 0.1536
5.45$\uparrow$ 4.30$\uparrow$ 4.83$\uparrow$ 3.14$\uparrow$ 6.40$\uparrow$ 4.84$\uparrow$ 5.56$\uparrow$ 6.10$\uparrow$ 6.25$\uparrow$ 4.89$\uparrow$ 4.79$\uparrow$ 3.99$\uparrow$
Yelp2018
Encoder MF 0.0429 0.0246 0.0284 0.0470 0.0365 0.0417 0.0352 0.0491 0.0305 0.0700 0.0394 0.0580
LightGCN 0.0527 0.0318 0.0354 0.0577 0.0452 0.0536 0.0430 0.0613 0.0374 0.0884 0.0464 0.0724
NGCF 0.0511 0.0307 0.0343 0.0555 0.0438 0.0523 0.0419 0.059 0.0364 0.0865 0.0453 0.0702
Negative DNS 0.0469 0.0279 0.0312 0.0509 0.0404 0.0481 0.0385 0.0541 0.0341 0.0804 0.0422 0.0647
PNS 0.0324 0.0196 0.0218 0.0351 0.0280 0.0335 0.0267 0.0374 0.0235 0.0555 0.0291 0.0448
HarS 0.0418 0.0241 0.0273 0.0456 0.0353 0.0407 0.0331 0.0474 0.0295 0.0678 0.0362 0.0559
2stS 0.0393 0.0234 0.0261 0.0429 0.0335 0.0397 0.0319 0.0454 0.0277 0.0654 0.0342 0.0536
Loss CML 0.0347 0.0221 0.0245 0.0397 0.0309 0.0386 0.0301 0.0398 0.0261 0.0658 0.0314 0.0501
SFCML 0.0552 0.0332 0.0370 0.0605 0.0469 0.0556 0.0448 0.0637 0.0384 0.0903 0.0477 0.0747
DPCML 0.0467 0.0273 0.0307 0.0506 0.0393 0.0459 0.0371 0.053 0.0328 0.076 0.0403 0.0626
DCL 0.0485 0.0285 0.0321 0.0528 0.0415 0.0485 0.0391 0.0557 0.0346 0.0804 0.0426 0.0656
HCL 0.054 0.0327 0.0365 0.0591 0.0456 0.0544 0.0436 0.0621 0.0377 0.0886 0.0464 0.0726
BUIR 0.0597 0.0364 0.0404 0.0655 0.0501 0.0603 0.0481 0.0688 0.041 0.0982 0.0512 0.0809
Proposed OneBP 0.0616 0.0366 0.041 0.0675 0.0521 0.0613 0.0494 0.0708 0.0428 0.0997 0.0527 0.0827
3.18$\uparrow$ 0.55$\uparrow$ 1.49$\uparrow$ 3.05$\uparrow$ 3.99$\uparrow$ 1.66$\uparrow$ 2.70$\uparrow$ 2.91$\uparrow$ 4.39$\uparrow$ 1.53$\uparrow$ 2.93$\uparrow$ 2.22$\uparrow$
@@
    The value ahead $\uparrow$ measures the improvement percentage compared to the optimal value among other baselines.
@@
Table 3. Comparison of computation time per epoch in model training (s).
Dataset Encoder Negative Sampling Loss Ours
MF LightGCN NGCF DNS PNS HarS 2stS CML SFCML DPCML DCL HCL BUIR OneBP
100k 1.19 2.17 3.21 1.83 12.09 1.59 13.45 1.45 2.61 1.55 1.24 1.26 1.57 1.11 6.72$\downarrow$
1M 14.28 47.34 62.35 19.36 254.74 18.24 276.35 16.39 57.02 16.89 14.24 14.62 16.58 14.21 0.49$\downarrow$
Gowalla 20.91 273.50 157.74 37.05 3728.37 31.52 3125.77 91.11 1842.02 92.13 21.12 20.98 30.17 20.12 3.78$\downarrow$
Yelp2018 44.02 1155.24 502.13 73.58 5526.33 54.75 4521.21 143.08 3482.52 144.78 44.29 44.5s 64.80 36.33 17.50$\downarrow$
@@ 5.4. Hyper-parameter Study
Figure 6. Performance vs. number of negative samples.

### v3

@@ 2. Two Tower Recommendation Models
$\mathbf{u}_{i}=UserEncoder(u_{i}),\mathbf{v}_{j}=ItemEncoder(v_{j}).$
@@ 5.1. Experiment Settings
[…]
Parameter Settings: To make fair comparisons, we use the same number of negative samples for all algorithms. More detailed hyperparameter settings are documented in the README file of the code. The codes is implemented with Pytorch  [^4] Codes are released at: https://anonymous.4open.science/r/OneBP.
@@ 5.2. Overall results
Table 2 also compares the computation time per epoch on the four datasets, where our OneBP requires the least ones. The savings of computation time become more significant in larger datasets. This is not unexpected. As our solution does not require advanced neural encoding, sophisticated negative sampling and complicated loss computing, its computation complexity for model training can be greatly reduced. Further computation savings can be attributed to that our OneBP does not need to compute loss gradients for updating users’ encodings.
@@
Dataset Type Method P@5 R@5 F1@5 NDCG@5 P@10 R@10 F1@10 NDCG@10 P@20 R@20 F1@20 NDCG@20 time(s)
MovieLens
100k
Encoder MF 0.4155 0.1434 0.1850 0.4430 0.3496 0.2308 0.2328 0.4181 0.2824 0.3529 0.2606 0.4183 1.19
LightGCN 0.4263 0.1458 0.1881 0.4531 0.3587 0.2313 0.2351 0.4269 0.2877 0.3474 0.2609 0.4236 2.17
NGCF 0.4176 0.1452 0.1871 0.4469 0.3582 0.2356 0.238 0.4264 0.2875 0.3597 0.2652 0.4248 3.21
Negative DNS 0.4214 0.1460 0.1882 0.4501 0.3531 0.2317 0.2342 0.4226 0.2836 0.3544 0.2614 0.4228 1.83
PNS 0.3979 0.1338 0.1742 0.4231 0.3415 0.218 0.223 0.4026 0.2741 0.3373 0.2511 0.4016 12.09
HarS 0.4227 0.1455 0.1878 0.4516 0.3527 0.2312 0.2342 0.4239 0.2852 0.3557 0.2624 0.4247 1.59
2stS 0.3718 0.1217 0.1597 0.3929 0.3188 0.2027 0.2085 0.3715 0.2574 0.3184 0.2365 0.3717 13.45
Loss CML 0.3796 0.1174 0.1562 0.3963 0.3314 0.2022 0.2109 0.3793 0.272 0.3243 0.2458 0.3811 1.45
SFCML 0.4174 0.1432 0.1846 0.4429 0.3544 0.2292 0.2321 0.4202 0.2814 0.3389 0.2542 0.4147 2.61
DPCML 0.3869 0.1337 0.1717 0.4115 0.3207 0.2108 0.213 0.3836 0.2647 0.3287 0.2435 0.3833 1.55
DCL 0.4291 0.1501 0.1928 0.4587 0.3608 0.2384 0.2402 0.4319 0.2845 0.3542 0.2621 0.4257 1.24
HCL 0.4263 0.1499 0.1924 0.4576 0.3576 0.2365 0.2389 0.4295 0.2814 0.3513 0.2603 0.4229 1.26
BUIR 0.4146 0.1387 0.1804 0.4373 0.3491 0.2258 0.2297 0.4121 0.2743 0.3445 0.2529 0.407 1.57
Proposed OneBP 0.453 0.1559 0.2015 0.4822 0.3811 0.2494 0.2529 0.4535 0.3032 0.373 0.2776 0.4501 1.11
5.57$\uparrow$ 3.86$\uparrow$ 4.51$\uparrow$ 5.12$\uparrow$ 5.63$\uparrow$ 4.61$\uparrow$ 5.29$\uparrow$ 5.00$\uparrow$ 5.39$\uparrow$ 3.70$\uparrow$ 4.68$\uparrow$ 5.73$\uparrow$ 6.72$\downarrow$
MovieLens
1M
Encoder MF 0.4046 0.0962 0.1355 0.4234 0.3510 0.1584 0.1816 0.3930 0.2922 0.2477 0.2166 0.3780 14.28
LightGCN 0.4121 0.0955 0.1350 0.4327 0.3537 0.1584 0.1792 0.3981 0.2921 0.2401 0.2132 0.3791 47.34
NGCF 0.4192 0.1014 0.1422 0.4397 0.3608 0.1643 0.1879 0.4059 0.2983 0.2560 0.2234 0.3888 62.35
Negative DNS 0.4080 0.0975 0.1372 0.4276 0.3558 0.1606 0.1841 0.3980 0.2948 0.2504 0.2190 0.3819 19.36
PNS 0.3931 0.0950 0.1334 0.4129 0.3362 0.1542 0.1759 0.3802 0.2773 0.2386 0.2079 0.3639 254.74
HarS 0.3453 0.0719 0.1042 0.3687 0.2770 0.1086 0.1293 0.3201 0.2049 0.1544 0.1424 0.2781 18.24
2stS 0.3772 0.0930 0.1295 0.3978 0.3261 0.1507 0.1704 0.3679 0.2689 0.2362 0.2029 0.3521 276.35
Loss CML 0.3425 0.0753 0.1077 0.3554 0.3047 0.1277 0.1497 0.3341 0.258 0.2055 0.1847 0.3226 16.39
SFCML 0.3968 0.0959 0.1348 0.4136 0.3481 0.1573 0.1808 0.3875 0.2912 0.243 0.2154 0.3741 57.02
DPCML 0.3883 0.088 0.125 0.4079 0.3359 0.1424 0.1665 0.3755 0.2758 0.2198 0.1983 0.3538 16.89
DCL 0.4133 0.0953 0.1350 0.4355 0.3512 0.152 0.177 0.3967 0.2867 0.2328 0.2084 0.3732 14.24
HCL 0.4125 0.0957 0.1355 0.4345 0.3515 0.1528 0.1776 0.3972 0.2867 0.2336 0.2086 0.3736 14.62
BUIR 0.3875 0.0942 0.1318 0.4089 0.3314 0.152 0.1728 0.376 0.2714 0.2332 0.2023 0.3581 16.58
Proposed OneBP 0.4300 0.1071 0.1487 0.4512 0.3709 0.1746 0.1966 0.4182 0.3069 0.2690 0.2316 0.4036 14.21
2.58$\uparrow$ 5.62$\uparrow$ 4.57$\uparrow$ 2.62$\uparrow$ 2.80$\uparrow$ 6.27$\uparrow$ 4.63$\uparrow$ 3.03$\uparrow$ 2.88$\uparrow$ 5.08$\uparrow$ 3.67$\uparrow$ 3.81$\uparrow$ 0.49$\downarrow$
Gowalla Encoder MF 0.0739 0.0757 0.0621 0.1016 0.0560 0.1122 0.0623 0.1076 0.0422 0.1650 0.0587 0.1230 20.91
LightGCN 0.0845 0.0888 0.0725 0.1166 0.0631 0.1297 0.0713 0.1232 0.0465 0.1864 0.0642 0.1394 273.5
NGCF 0.0768 0.0808 0.0661 0.1051 0.0585 0.121 0.0664 0.1127 0.044 0.1769 0.0608 0.1291 157.74
Negative DNS 0.0737 0.0779 0.0633 0.1013 0.0563 0.1162 0.0636 0.1085 0.0421 0.1696 0.0580 0.1240 37.05
PNS 0.0586 0.0611 0.0502 0.0811 0.0439 0.0894 0.0494 0.0855 0.0323 0.1298 0.0445 0.0968 3728.37
HarS 0.0719 0.0742 0.0609 0.0989 0.0544 0.1098 0.0607 0.1048 0.0411 0.16 0.0558 0.1189 31.52
2stS 0.0609 0.0599 0.0501 0.0832 0.0461 0.0883 0.0502 0.0871 0.0345 0.1285 0.0462 0.0982 3125.77
Loss CML 0.0501 0.0603 0.0472 0.0705 0.0384 0.0901 0.0466 0.0782 0.0284 0.1285 0.0412 0.0902 91.11
SFCML 0.0899 0.0931 0.0766 0.1243 0.0672 0.1365 0.0756 0.1278 0.0496 0.1963 0.0689 0.1477 1842.02
DPCML 0.0782 0.0759 0.0642 0.1042 0.059 0.1124 0.0645 0.1092 0.0432 0.1632 0.0584 0.1225 92.13
DCL 0.0761 0.0776 0.0642 0.1036 0.0575 0.1149 0.064 0.1093 0.0424 0.1651 0.0578 0.1236 21.12
HCL 0.0873 0.0885 0.0731 0.118 0.0653 0.1286 0.0723 0.1236 0.0477 0.1846 0.0649 0.1392 20.98
BUIR 0.0854 0.0872 0.072 0.1166 0.0651 0.1306 0.0725 0.1241 0.0484 0.1898 0.0661 0.1411 30.17
Proposed OneBP 0.0948 0.0971 0.0803 0.1282 0.0715 0.1431 0.0798 0.1356 0.0527 0.2059 0.0722 0.1536 20.12
5.45$\uparrow$ 4.30$\uparrow$ 4.83$\uparrow$ 3.14$\uparrow$ 6.40$\uparrow$ 4.84$\uparrow$ 5.56$\uparrow$ 6.10$\uparrow$ 6.25$\uparrow$ 4.89$\uparrow$ 4.79$\uparrow$ 3.99$\uparrow$ 3.78$\downarrow$
Yelp2018 Encoder MF 0.0429 0.0246 0.0284 0.0470 0.0365 0.0417 0.0352 0.0491 0.0305 0.0700 0.0394 0.0580 44.02
LightGCN 0.0527 0.0318 0.0354 0.0577 0.0452 0.0536 0.0430 0.0613 0.0374 0.0884 0.0464 0.0724 1155.24
NGCF 0.0511 0.0307 0.0343 0.0555 0.0438 0.0523 0.0419 0.059 0.0364 0.0865 0.0453 0.0702 502.13
Negative DNS 0.0469 0.0279 0.0312 0.0509 0.0404 0.0481 0.0385 0.0541 0.0341 0.0804 0.0422 0.0647 73.58
PNS 0.0324 0.0196 0.0218 0.0351 0.0280 0.0335 0.0267 0.0374 0.0235 0.0555 0.0291 0.0448 5526.33
HarS 0.0418 0.0241 0.0273 0.0456 0.0353 0.0407 0.0331 0.0474 0.0295 0.0678 0.0362 0.0559 54.75
2stS 0.0393 0.0234 0.0261 0.0429 0.0335 0.0397 0.0319 0.0454 0.0277 0.0654 0.0342 0.0536 4521.21
Loss CML 0.0347 0.0221 0.0245 0.0397 0.0309 0.0386 0.0301 0.0398 0.0261 0.0658 0.0314 0.0501 143.08
SFCML 0.0552 0.0332 0.0370 0.0605 0.0469 0.0556 0.0448 0.0637 0.0384 0.0903 0.0477 0.0747 3482.52
DPCML 0.0467 0.0273 0.0307 0.0506 0.0393 0.0459 0.0371 0.053 0.0328 0.076 0.0403 0.0626 144.7
DCL 0.0485 0.0285 0.0321 0.0528 0.0415 0.0485 0.0391 0.0557 0.0346 0.0804 0.0426 0.0656 44.29
HCL 0.054 0.0327 0.0365 0.0591 0.0456 0.0544 0.0436 0.0621 0.0377 0.0886 0.0464 0.0726 44.5
BUIR 0.0597 0.0364 0.0404 0.0655 0.0501 0.0603 0.0481 0.0688 0.041 0.0982 0.0512 0.0809 64.8
Proposed OneBP 0.0616 0.0366 0.041 0.0675 0.0521 0.0613 0.0494 0.0708 0.0428 0.0997 0.0527 0.0827 36.33
3.18$\uparrow$ 0.55$\uparrow$ 1.49$\uparrow$ 3.05$\uparrow$ 3.99$\uparrow$ 1.66$\uparrow$ 2.70$\uparrow$ 2.91$\uparrow$ 4.39$\uparrow$ 1.53$\uparrow$ 2.93$\uparrow$ 2.22$\uparrow$ 17.5$\downarrow$
@@
    The value ahead $\uparrow$ and $\downarrow$ measures the improvement and degradation percentage compared to the optimal value among other baselines.
@@ 5.4. Hyper-parameter Study
Figure 5. Hyper-parameter Study for $N_{s}$ and $\beta$